For Christmas break, you decide to head to Saginaw, Texas. The distance you expect to travel is 480 miles, and your car gets 26.5 miles per gallon. If the cost of gasoline is $2.82 per gallon, how much money will you need for gas to get to Saginaw?
Answer #1

here,

the distance traveled , s = 480 mi

the gasoline used , V = s /26.5 gallon

V = 480 /26.5 gallons

V = 18.11 gallons

the money needed , C = V * 2.82 dollars

C = 18.11 * 2.82 dollars

C = 51.1 dollars

the money needed is 51.1 dollars